express 0.0272727... in to quotient form solve |x-1|=0.1
Answer by oscargut(2103)

|x-1|=0.1
x-1 = 0.1 or x-1 = -0.1
x = 1.1 or x = 0.9
If you have more questions i can see them if you contact me at: mthman@gmail.com
Thanks